OpenStudy (anonymous):

HELP PLEASEEEEEEEEEEE I'm MELTINGGGGGGGGGGG If Susan invests $1000 in an account that yields 5% annual, compounding quarterly, then how much does she have after 6 years? Why do I divide 5/4 instead of 1.05/4?

OpenStudy (anonymous):

um i think that for this you will have to multiply .05 and 6 and then multipy 1000 by 1.** the decimal with be the .05 * 6 sorry if im wrong im just trying to help out so dont hate

OpenStudy (jdoe0001):

if you're melting.... I'd suggest some ice cubes, you may want to get some kool-aid with it too

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$1000\\ r=rate\to 5\%\to \frac{5}{100}\to 0.05\\ n=\textit{times it compounds per year, quarterly means}\to 4\\ t=years\to 6 \\ \quad \\ A=1000\left(1+\frac{0.05}{4}\right)^{4\cdot 6}\)